OpenStudy (swaggyswags):

Let f(x) = 2x − 6. Solve f−1(x) when x = 2. −5 4 −2 7

OpenStudy (swaggyswags):

@mathstudent55

OpenStudy (swaggyswags):

Let f(x) = 9x − 2 and g(x) = −x + 3. Find f(g(x)). −9x − 2 −9x + 5 −9x + 25 −9x + 27

OpenStudy (p0sitr0n):

FOr the first, the inverse is f(x)=x/2+3. Plugging in x=2 gives f-1(2)=4. For the second, f(g(x))=9(-x+3)-2=-9x+27-2=-9x+25.

OpenStudy (karim728):